# Multi-Modal Semantic Parsing for the Automatic Interpretation of Tombstone Inscriptions

### v1

@@ Abstract.
Tombstones are historically and culturally rich artifacts, encapsulating individual lives, community memory, historical narratives and artistic expression. Yet, many tombstones today face significant preservation challenges, including physical erosion, vandalism, environmental degradation, and political shifts. In this paper, we introduce a novel multi-modal framework for tombstones digitization, aiming to improve the interpretation, organization and retrieval of tombstone content. Our approach leverages vision-language models (VLMs) to translate tombstone images into structured Tombstone Meaning Representations (TMRs), capturing both image and text information. To further enrich semantic parsing, we incorporate retrieval-augmented generation (RAG) for integrate externally dependent elements such as toponyms, occupation codes, and ontological concepts. Compared to traditional OCR-based pipelines, our method improves parsing accuracy from an F1 score of 36.1 to 89.5. We additionally evaluate the model’s robustness across diverse linguistic and cultural inscriptions, and simulate physical degradation through image fusion to assess performance under noisy or damaged conditions. Our work represents the first attempt to formalize tombstone understanding using large vision-language models, presenting implications for heritage preservation.
@@
isbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Tombstone inscriptions give us rich historical and cultural insight. Therefore, cemeteries have sparked the interest of many different types of researchers, including anthropologists, archaeologists, genealogists, historians, philologists, sociologists, and theologists (Saller and Shaw, 1984; Eckert, 1993; Veit and Nonestied, 2008; Long, 2016; Bos, 2022). Besides the names of the deceased and dates and location of birth and death, tombstones often also contain information about family relationships, occupations, epitaphs, biblical quotes, poetry, and more. However, the intricate nature of these tombstones makes it challenging to interpret, organize and retrieve the valuable information they contain. Moreover, many tombstones are at risk due to physical erosion, vandalism, environmental degradation, and political shifts (Matthias, 1967; Streiter et al., 2007a). Digital preservation is therefore essential—not only to safeguard these irreplaceable cultural records but also to facilitate storage and search in structured databases.
With the ultimate aim of creating a large database with structured information of tombstone inscriptions, we investigate the possibility of automatically parsing tombstone images into structured data suitable for storage, in order to assist data curators and provide cemetery researchers with a powerful tool for searching into databases of digitized tombstones. The overall idea can be seen in Figure 1, where the input is an image (photograph) of a tombstone, and the output is an entry in a relational database. [^1] All personal names shown in this paper have been blurred to protect individual privacy. Cemetery researchers are then able to systematically search tombstones on dates, toponyms, gender, jobs, relationships, and other features.
@@
  • How do models handle challenging tombstones, such as those with linguistically complex inscriptions or physical defects?
Following a review of background and related work on semantic parsing and tombstone meaning representations (Section 2), we introduce the T2M framework (Section 3). We then present a series of experiments (Section 4) that evaluate VLMs and RAG under zero-shot, few-shot, and fine-tuned settings, including three strategies for integrating retrieval components. Our results (Section 5) show that fine-tuned VLMs achieve state-of-the-art performance in tombstone parsing, with RAG further boosting accuracy by incorporating external domain knowledge. In Section 6, we assess model performance on linguistically challenging inscriptions across five dimensions, and evaluate robustness under simulated physical degradation using noise-infused tombstone images.
@@ 2.1. Automated Formal Semantic Parsing
Formal meaning representations play a crucial role in advancing natural language understanding by providing structured, abstract depictions of sentence meaning. Abstract Meaning Representation (AMR) (Banarescu et al., 2013) utilizes directed acyclic graphs to encapsulate the core semantic relationships within a sentence. Discourse Representation Structure (DRS) (Kamp and Reyle, 1993; Bos et al., 2017; Bos, 2023) is designed to capture more context-dependent phenomena such as anaphora, temporal expressions and quantification across extended discourse. There are many other semantic formalisms, such as BMR (Martínez Lorenzo et al., 2022), PTG (Hajič et al., 2012), and EDS (Oepen and Lønning, 2006), focusing on different semantic phenomena.
Semantic parsing is a cornerstone task in natural language processing that focuses on converting natural language text into structured formal meaning representations. This transformation involves mapping complex linguistic input into abstract structures—such as AMR, DRS, and others. Traditionally, this process has been tackled using rule-based systems (Woods, 1973; Hendrix et al., 1977; Templeton and Burger, 1983) as well as neural models (Barzdins and Gosko, 2016; van Noord and Bos, 2017; van Noord et al., 2018, 2020; Bevilacqua et al., 2021; Wang et al., 2023; Ettinger et al., 2023; Liu, 2024a, b; Yang et al., 2024; Zhang et al., 2024b, 2025), primarily with textual input.
Recent advancements have extended semantic parsing into the multimodal domain, integrating non-linguistic modalities such as vision, speech or gesture. For instance, Abdelsalam et al. (2022) introduced visual semantic parsing, translating scene graphs extracted from images into AMR-like representations. Brutti et al. (2022) proposed gesture meaning representation, capturing the semantic content of hand gestures. Multimodal AMR parsing has also been explored, combining visual and textual inputs for enhanced grounding (Bonial et al., 2023; Lai et al., 2024). There are also many works in the field of Robotics. Thomason et al. (2020) used multi-modal learning (robotic sensors) to improve perceptual concept parsing for human-robot interaction. Farazi et al. (2020) introduced a semantic relationship parser for visual question answering, which performs multi-modal semantic parsing by generating semantic feature vectors from images and questions.
@@ 3. Methodology
Figure 2. Overview of the Tomb2Meaning (T2M) framework. The three retrieval-augmented generation (RAG) strategies are: Retrieval-in-the-Beginning Augmented Generation (RibAG), Retrieval-in-the-Middle Augmented Generation (RimAG), and Retrieval-in-the-End Augmented Generation (RieAG). We denote two modes of model usage: direct prompting (indicated by the snowflake symbol) and fine-tuning (indicated by the flame symbol). Prompt templates and training settings are in Appendix B.
@@ 3.1. Baselines
To establish a low bound of performance, we introduce the deterministic, input-agnostic baseline parser called Deterministic Baseline (May, 2016). It has been created by computing the average Smatch score for each TMR across training pairs. The TMR  [^2] This deterministic TMR can be seen in Figure 8 in Appendix A. with the highest average score is selected as a constant output, representing the minimal performance achievable by any well-designed system.
@@ 3.2. The Tomb2Meaning framework
    Query Construction: Extracted entities $E(\cdot)$, derived from either the input image or an initially generated TMR, are used to construct retrieval queries, which are then sent to the APIs of GeoNames, HISCO, and WordNet [^3] Details of search are in Appendix B..
@@
    Prompt/Instruction Reconstruction: Once the external retrieval information is obtained, the prompts (or fine-tuning instructions) are reconstructed to incorporate this supplementary data, as detailed in Appendix B.
@@ 4.1. Dataset
While most inscriptions are in Dutch and rendered in serif fonts, the dataset also includes samples in English, French, German, Spanish, Italian, Indonesian, and Greek. Gothic script is another frequently occurring typeface. A small number of inscriptions exhibit multilingual mixing. Example images and corresponding annotations are provided in Appendix A.
For experimental purposes, we have randomly divided these images into equal-sized training and testing sets (50:50 split), ensuring a sufficiently large sample for testing.
@@ 4.3. Models and Settings
We conducted two experiments on the base framework: one using a frozen VLM via prompting and the other by fine-tuning the VLM. In the prompting experiment, we compared model performance using few-shot examples that were randomly selected from the train set. For the fine-tuning, we applied parameter-efficient fine-tuning (PEFT), specifically using LoRA(Hu et al., 2022).
All experiments are conducted on a standardized hardware setup, with detailed configurations provided in Appendix B to support replication and future comparative studies. [^5] The code will be released after the anonymous review period.
@@ 5.1. Few-shot Prompting
Figure 3. Smatch scores for five models (solid lines) and three baselines (dashed lines) under few-shot prompting, with the number of shots ranging from 0 to 6. Numerical results are provided in Appendix C.
Table 2. Performance of the automated tombstone parsing framework (T2M) with and without retrieval augmented generation methods for various models, compared to three baselines. Here, Smatch is the overall F1-score, F1-X represents the F1-score for instances of X, and IFR is the ill-formed frequency rate of the parser output.
@@ 6.2. Challenges of Tombstone Images
In real-world scenarios, tombstones may exhibit defects such as cracks, erosion, or other damage that affect legibility. While our annotated dataset mainly includes clear inscriptions, we simulate defective tombstones to evaluate robustness. We employ an image fusion technique via alpha blending (see Figure 6), where original images are blended with artificially generated damage. Noise levels are parameterized into low, medium, and high settings (detailed in Appendix D).
@@ 7. Conclusion
Appendix A Examples of Tombstone&Meaning Representation
In Figure 8, we provide three examples of tombstones and their TMRs in Penman notation and graph notation.
Figure 8. Three examples of tombstones and meaning representations in Penman notation and graph notation. t00294 is the deterministic output in our baseline.
Appendix B Experiment Details
In this section, we provide the details of our experiments, including the prompts we used, the setting of fine tuning, the numeric results.
Prompts for Tombstone Parsing
Zero-shot Prompt ¡Image¿Generate a meaning representation in PENMAN format for this image of a tombstone. Dont give any other text or explanations.
few-shot Prompt ¡Image1¿ ¡TMR1¿ ¡Image2¿ ¡TMR2¿ ¡Image¿The above are examples of meaning representation in PENMAN format for a tombstone for the tombstone images. Please generate a meaning representation in PENMAN format for this image of a tombstone. Dont give any other text or explanations.
few-shot Prompt + RAG ¡Image1¿ ¡TMR1¿ ¡Image2¿ ¡TMR2¿ ¡Image¿The above are examples of meaning representation in PENMAN format for a tombstone for the tombstone images. Please generate a meaning representation in PENMAN format for this image of a tombstone. he GeoCode you may use are:place1: code1, place2: code2, …. The HISCO code you may use are: occupation1: code1, occupation2: code2, …. The WordNet sysnet you may use are: concept1: definition1, concept2: …. Dont give any other text or explanations.
Prompts for Tombstone OCR
Prompt for OCR ¡Image¿Below is an image of a tombstone. Please perform OCR on the image and return the its inscriptions, without adding any extra commentary or interpretation.
Instruction for Fine-Tuning
T2M instruction ¡Image¿Generate a meaning representation in PENMAN format for this image of a tombstone.
T2M+RAG instruction ¡Image¿Generate a meaning representation in PENMAN format for this image of a tombstone. The GeoCode you may use are:{place1: code1, place2: code2, …}. The HISCO code you may use are: {occupation1: code1, occupation2: code2, …}. The WordNet sysnet you may use are: {concept1: definition1, concept2: …}.
Search Engine
The API of search:
Fine-tuning settings
Table 3 provide the basic details of the experiments.
Table 3. Configurations for large language models Fine-Tuning and Inference.
Category Details Category Details
Stage SFT/inference Precision fp16
Fine-tuning LoRA Batch Size 1
Cutoff Length 1024 GPU Number 1
Learning Rate $10^{-4}$ GPU H100
Epochs 10 lr scheduler cosine
Appendix C Performances of Few-shot Prompting
Table 4. Numeric performance of Few-shot prompting.
Model Size Prompt T2M T2M+RibAG T2M+RimAG T2M+RieAG
Qwen2.5-VL 3B zero-shot 0.00 0.00 0.00 0.00
one-shot 14.52 18.12 20.95 23.34
two-shot 16.94 18.77 20.81 19.58
three-shot 22.51 26.68 27.27 28.90
four-shot 23.82 24.14 27.56 27.39
five-shot 25.44 25.23 28.04 28.77
six-shot 24.98 25.47 27.35 28.66
7B zero-shot 0.00 0.00 0.00 0.00
one-shot 56.81 59.86 62.12 58.44
two-shot 60.79 65.34 66.59 68.11
three-shot 65.42 68.43 70.25 69.97
four-shot 65.93 65.90 69.41 68.02
five-shot 66.12 69.15 70.68 70.50
six-shot 67.88 68.76 70.33 69.84
72B zero-shot 0.00 0.00 0.00 0.00
one-shot 42.98 47.92 48.57 49.18
two-shot 63.88 67.85 69.23 68.62
three-shot 66.22 68.34 71.90 70.49
four-shot 66.50 70.81 72.65 69.37
five-shot 67.10 69.43 71.99 70.15
six-shot 68.64 69.70 71.48 70.27
Llava-v1.6-mistral 7B zero-shot 0.00 0.00 0.00 0.00
one-shot 36.58 38.05 39.47 37.38
two-shot 42.19 45.29 48.53 47.61
three-shot 50.21 56.92 58.15 57.68
four-shot 55.45 59.81 60.64 60.07
five-shot 58.84 60.55 63.97 61.24
six-shot 59.12 61.76 64.20 62.45
Llama-3.2-vision-11B 11B zero-shot 0.00 0.00 0.00 0.00
one-shot 38.97 40.68 45.35 43.29
two-shot 40.88 43.97 46.88 44.54
three-shot 54.36 56.43 58.77 57.96
four-shot 56.85 59.64 60.29 59.14
five-shot 59.48 61.89 63.76 63.40
six-shot 59.15 60.22 61.15 60.15
Table 4 shows the numeric results of few-shot prompting on tombstone parsing, as a supplement for the line chart in main text.
Appendix D Image Fusion
Each damage image undergoes preprocessing to remove a fixed background color and isolate damage features with an alpha channel. Subsequently, we generate complex masks to give natural shapes to the damage artifacts. These masks, randomly shaped as clouds, slender stripes, or irregular ellipses, are created using operations like noise addition, Gaussian blurring, and morphological transformations.
The damage patches are resized proportionally to the tombstone images and blended through per-pixel alpha blending, controlling transparency within a predefined range for subtle integration. Multiple overlays with varied scaling and blending weights produce augmented images that effectively simulate realistic deterioration.

### v3

# Multi-Modal Semantic Parsing for the Interpretation of Tombstone Inscriptions

@@ Abstract.
Tombstones are historically and culturally rich artifacts, encapsulating individual lives, community memory, historical narratives and artistic expression. Yet, many tombstones today face significant preservation challenges, including physical erosion, vandalism, environmental degradation, and political shifts. In this paper, we introduce a novel multi-modal framework for tombstones digitization, aiming to improve the interpretation, organization and retrieval of tombstone content. Our approach leverages vision-language models (VLMs) to translate tombstone images into structured Tombstone Meaning Representations (TMRs), capturing both image and text information. To further enrich semantic parsing, we incorporate retrieval-augmented generation (RAG) for integrate externally dependent elements such as toponyms, occupation codes, and ontological concepts. Compared to traditional OCR-based pipelines, our method improves parsing accuracy from an F1 score of 36.1 to 89.5. Furthermore, we evaluate the model’s robustness across diverse linguistic and cultural inscriptions, and simulate physical degradation through image fusion to assess performance under noisy or damaged conditions. Our work represents the first attempt to formalize tombstone understanding using large vision-language models, presenting implications for heritage preservation. The code and supplementary materials are available at: https://github.com/LastDance500/Tombstone-Parsing.
@@
journalyear: 2025copyright: acmcopyrightconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755444isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Natural language generationccs: Computing methodologies Image representations
@@ 1. Introduction
Tombstone inscriptions give us rich historical and cultural insight. Therefore, cemeteries have sparked the interest of many different types of researchers, including anthropologists, archaeologists, genealogists, historians, philologists, sociologists, and theologists (Saller and Shaw, 1984; Eckert, 1993; Carmack, 2002; Veit and Nonestied, 2008; Long, 2016; Bos, 2022; Chen et al., 2024). Besides the names of the deceased and dates and location of birth and death, tombstones often also contain information about family relationships, occupations, epitaphs, biblical quotes, poetry, and more. However, the intricate nature of these tombstones makes it challenging to interpret, organize and retrieve the valuable information they contain. Moreover, many tombstones are at risk due to physical erosion, vandalism, environmental degradation, and political shifts (Matthias, 1967; Streiter et al., 2007a). Digital preservation is therefore essential—not only to safeguard these irreplaceable cultural records but also to facilitate storage and search in structured databases.
With the ultimate aim of creating a large database with structured information of tombstone inscriptions, we investigate the possibility of automatically parsing tombstone images into structured data suitable for storage, in order to assist data curators and provide cemetery researchers with a powerful tool for searching into databases of digitized tombstones. The overall idea can be seen in Figure 1, where the input is an image (photograph) of a tombstone, and the output is an entry in a relational database. [^1] Names of people on tombstones shown in this paper have been masked. Our aim is to stimulate research in developing automatic tools for cemetery research rather than distributing private information online. Cemetery researchers are then able to systematically search tombstones on dates, toponyms, gender, jobs, relationships, and other features.
@@
  • How do models handle challenging stones, such as those with linguistically complex inscriptions or physical defects?
Following a review of background and related work on semantic parsing and tombstone meaning representations (Section 2), we introduce the T2M framework (Section 3). We then present the experiments (Section 4) that evaluate VLMs and RAG under few-shot and fine-tuned settings, including three strategies for integrating retrieval components. Our results (Section 5) show that fine-tuned VLMs achieve the state-of-the-art performance in tombstone parsing, with RAG further boosting accuracy by incorporating external domain knowledge. In Section 6, we assess the models on linguistically challenging inscriptions across five dimensions, and evaluate robustness under simulated physical degradation.
@@ 2.1. Automated Formal Semantic Parsing
Formal meaning representations play a crucial role in advancing natural language understanding (NLP) by providing structured, abstract depictions of sentence meaning. Abstract Meaning Representation (AMR) (Banarescu et al., 2013) utilizes directed acyclic graphs to encapsulate the core semantic relationships within a sentence. Discourse Representation Structure (DRS) (Kamp and Reyle, 1993; Bos et al., 2017; Bos, 2023) is designed to capture more context-dependent phenomena such as anaphora, temporal expressions and quantification across extended discourse. There are many other semantic formalisms, such as BMR (Mart´ınez Lorenzo et al., 2022), PTG (Hajič et al., 2012), and EDS (Oepen and Lønning, 2006), focusing on different semantic phenomena.
Semantic parsing is a cornerstone task in NLP that focuses on converting natural language text into structured formal meaning representations. This transformation involves mapping complex linguistic input into abstract structures—such as AMR, DRS, and others. Traditionally, this process has been tackled using rule-based systems (Woods, 1973; Hendrix et al., 1977; Templeton and Burger, 1983) as well as neural models (Barzdins and Gosko, 2016; van Noord and Bos, 2017; van Noord et al., 2018, 2020; Bevilacqua et al., 2021; Wang et al., 2023; Ettinger et al., 2023; Liu, 2024a, b; Yang et al., 2024; Zhang et al., 2024b, 2025), primarily with textual input.
Recent advancements have extended semantic parsing into the multimodal domain, integrating non-linguistic modalities such as vision, speech or gesture. For instance, Abdelsalam et al. (2022) introduced visual semantic parsing, translating scene graphs extracted from images into AMR-like representations. Brutti et al. (2022) proposed gesture meaning representation, capturing the semantic content of hand gestures. Multimodal AMR parsing has also been explored, combining visual and textual inputs for enhanced grounding (Bonial et al., 2023; Lai et al., 2024). There are also many works in the field of Robotics. Thomason et al. (2020) used multi-modal learning to improve perceptual concept parsing for human-robot interaction. Farazi et al. (2020) introduced a semantic relationship parser for visual question answering, which performs multi-modal semantic parsing by generating semantic feature vectors from images and questions.
@@ 3. Methodology
Figure 2. Overview of the Tomb2Meaning (T2M) framework. The three retrieval-augmented generation (RAG) strategies are: Retrieval-in-the-Beginning Augmented Generation (RibAG), Retrieval-in-the-Middle Augmented Generation (RimAG), and Retrieval-in-the-End Augmented Generation (RieAG). We denote two modes of model usage: direct prompting (indicated by the snowflake symbol) and fine-tuning (indicated by the flame symbol). Prompt templates and training settings are in supplemental materials.
@@ 3.1. Baselines
To establish a low bound of performance, we introduce the deterministic, input-agnostic baseline parser called Deterministic Baseline (May, 2016). It has been created by computing the average Smatch score for each TMR across training pairs. The TMR [^2] This deterministic TMR can be seen in the supplemental materials. with the highest average score is selected as a constant output, representing the minimal performance achievable by any well-designed system.
@@ 3.2. The Tomb2Meaning framework
    Query Construction: Extracted entities $E(\cdot)$, derived from either the input image or an initially generated TMR, are used to construct retrieval queries, which are then sent to the APIs of GeoNames, HISCO, and WordNet.
@@
    Prompt/Instruction Reconstruction: Once the external retrieval information is obtained, the prompts (or fine-tuning instructions) are reconstructed to incorporate this supplementary data.
@@ 4.1. Dataset
While most inscriptions are in Dutch and rendered in serif fonts, the dataset also includes samples in English, French, German, Spanish, Italian, Indonesian, and Greek. Gothic script is another frequently occurring typeface. A small number of inscriptions exhibit multilingual mixing.
For experimental purposes, we have randomly divided these images into equal-sized training and testing sets (50:50 split), ensuring a sufficiently large sample for testing. Example images, annotations and dataset statistics are provided in our supplemental materials.
@@ 4.3. Models and Settings
We conducted two experiments on the base framework: one using a frozen VLM via prompting and the other by fine-tuning the VLM. In the prompting experiment, we compared model performance using few-shot examples that were randomly selected from the train set. For the fine-tuning, we applied parameter-efficient fine-tuning (PEFT), specifically using LoRA(Hu et al., 2022). All experiments are conducted on a standardized hardware setup, with detailed configurations provided in supplemental materials to support replication and future comparative studies.
@@ 5.1. Few-shot Prompting
Figure 3. Smatch scores for five models (solid lines) and three baselines (dashed lines) under few-shot prompting, with the number of shots ranging from 0 to 6.
Table 2. Performance of the automated tombstone parsing framework (T2M) with and without retrieval augmented generation methods for various models, compared to three baselines. Here, Smatch is the overall F1-score, F1-X represents the F1-score for instances of X, and IFR is the ill-formed frequency rate of the parser output. All numbers are presented as percentages (%).
@@ 6.2. Challenges of Tombstone Images
In real-world scenarios, tombstones may exhibit defects such as cracks, erosion, or other damage that affect legibility. While our annotated dataset mainly includes clear inscriptions, we simulate defective tombstones to evaluate robustness. We employ an image fusion technique via alpha blending (see Figure 6), where original images are blended with artificially generated damage. Noise levels are parameterized into low, medium, and high settings (detailed in our supplemental materials).